Sales . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . $20,000

Variable expenses. . . . . . . . . . . . . . . . . . . . . . . . . . . . 12,000

Contribution margin . . . . . . . . . . . . . . . . . . . . . . . . . . 8,000

Fixed expenses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6,000

Net operating income. . . . . . . . . . . . . . . . . . . . . . . . . $ 2,000

Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

1. What is the contribution margin per unit?

2. What is the contribution margin ratio?

3. What is the variable expense ratio?

7. If the variable cost per unit increases by $1, spending on advertising increases by $1,500, and unit sales increase by 250 units, what would be the net operating income?

8. What is the break-even point in unit sales?

9. What is the break-even point in dollar sales?

11. What is the margin of safety in dollars? What is the margin of safety percentage?

12. What is the degree of operating leverage?

13. Using the degree of operating leverage, what is the estimated percent increase in net operating income of a 5% increase in sales?
Answer #1

Hi

Pls see below:

1. Contribution margin per unit = Contribution/units = 8000/1000 = 8 per unit

2.Contribution margin ratio = [Contribution/Sales] x 100 = [8000/20000] x100 = 40%

3. Variable expense ratio = [Variable expense/Sales] x 100 = [12000/20000] x100= 60%

7. Net operating income

Particulars $
Revised sales revenue 25000
[1250 units x 20 selling price]**
Less : Revised variable cost 16250
(12+1) x 1250 units **
Revised contribution margin 8750
Less : Revised Fixed cost [ 16000+ 1500] 7500
Net operating income 1250
** Notes:
Revised units = 1000+250 =1250
Selling price = 20000/1000 units = 20
Variable cost = 12000/1000 units = 12
Revised variable cost = 12+1 = 13
Revised Fixed cost = 6000+1500 advertisement = 7500

8. Break even point (units) = Fixed cost/contribution per unit = 6000/8= 750 units

9.Break even point (in dollars) = [Fixed cost/contribution margin ratio] = 6000/40%= $ 15000

11. Margin of safety in dollars = [Profit/contribution margin ratio] = 2000/40% = $ 5000

Margin of safety % = [Margin of safety sales/total sales] x100 = [5000/20000]x100 = 25%

12. Operating leverage = [Contribution/Operating profit] = 8000/2000 = 4

13. % increase in net operating income = [Sales increase % x Operating leverage] = 5% x 4 = 20%